OpenStudy (anonymous):

what is the integral of (x)/(1+x^4)

OpenStudy (anonymous):

PLEASE HELP.

OpenStudy (jack1):

(x)/(1+x^4) i'm pretty sure that integrates into a trig function, i'm going to pen and paper it dude, brb k?

OpenStudy (jack1):

yep, that's right integral of 1/x^2 + 1 = arctan (x) ... +c \[\int\limits_{}^{} \frac{ 1 }{ x^2 + 1 } = \tan^{-1} (x) + c\]

OpenStudy (jack1):

so for your q: \[\int\limits_{}^{} \frac {x}{x^4 +1} dx\] so sub in x^2 as u and make du = xdx so \[\int\limits_{}^{} \frac {1}{u^2 +1} du\] \[\frac 12 tan^{-1} (u) +c \] and u = x^2 so... \[\frac 12 tan^{-1} (x^2) +c \]

OpenStudy (anonymous):

One mistake... u = x^2 du = 2xdx

OpenStudy (anonymous):

\[\int \frac {x}{x^4 +1} dx\]\[=\frac{1}{2}\int \frac {1}{u^2 +1} du\]

OpenStudy (jack1):

...why 2x...? top line is x @rolypoly , where r u getting the 2x from dude?

OpenStudy (amistre64):

x - x^5 + x^9 + ... ----------------- 1 + x^4 | x (x + x^5) --------- -x^5 (-x^5 -x^9) ----------- x^9 (-x^9 +x^13) ------------ -x^13 etc ... \[\int(x - x^5 + x^9 + ...)~dx=\frac12x^2 - \frac16x^6 + \frac1{10}x^{10} + ...\] \[\sum_{0}\frac{(-1)^{n}}{2+4n}x^{2+4n}\]

OpenStudy (anonymous):

thanks Jack1. Thing is we were asked to solve for this equation using "u" substitution, and I couldn't do that. You were asking why 2x right, well if you make x^2= tan$, then when you differentiate both sides, you'll have 2x=sec^2$. Not to sure if I was clear when explaining, maybe someone else has a better explanation.

OpenStudy (jack1):

no, thats clear enough, cheers guys

OpenStudy (anonymous):

sweet. :)